Westway Company is a provider of imaging technology products and services to the photographic, graphic communications, and health-care markets. A condensed 2014 income statement follows (in millions):

Sales

$9,326

Costs of goods sold

7,197

Gross margin

2,129

Other operating expenses

2,896

Loss from continuing operations

$ (767)

Assume that $1,700 million of the cost of goods sold is a fixed cost representing depreciation and other production costs that do not change with the volume of production. In addition, $2,800 million of the other operating expenses is fixed.

  1. Compute the total contribution margin for 2014 and the contribution-margin percentage. Explain why the contribution margin differs from the gross margin.
  2. Suppose that sales for Westway are predicted to decrease by 10 percent or increase by either 10, 20 or 30 percent in 2015 and that the cost behaviour was expected to continue in 2015 as it did in 2014. Compute the predicted operating income (loss) for 2015 for each scenario.
  3. What assumptions were necessary to compute the predicted 2015 operating income in requirement 2?
Answer #1

Answer 1(a):

Fixed Costs of goods sold = $1700

Variable cost of goods sold = 7197 - 1700 = $5497

Fixed other operating expenses = $2,800

Variable other operating cost = 2,896 - 2800 = $96

Answer 1(b):

Contribution margin differs from the gross margin since:

1. It does not factor in fixed production costs

2. It factors in variable non-manufacturing costs

Answer 2:

Predicted operating income (loss) for 2015 for each scenario (highlighted yellow in $ millions) are as below:

Answer 3:

Assumptions necessary to compute the predicted 2015 operating income in requirement 2:

1. The cost behavior may change beyond a range. It is assumed fixed cost will remain $4500 for sales range from $8393 million to $12124 million

2. Variable cost per unit may also change beyond a range. It is assumed that variable cost per unit remains same for sales range from $8393 million to $12124 million